OpenStudy (anonymous):

Part 1: If you were solving a system of equations and you came to a statement like 4 = 4, what do you know about the system? Part 2: Solve the following system and show all of your work. 3y − 2x = 11 y + 2x = 9

OpenStudy (ivettef365):

you can add to solve for y as follows: 3y -2x = 11 y +2x = 9 ----------------

OpenStudy (anonymous):

1. when the results of an equation come out equal, like 4 = 4, that means there is infinite solutions. 2. y + 2x = 9 --> y = 9 - 2x now just sub 9 - 2x in for y in the 2nd equation, to find x. Once x is found,sub that back into the 1st equation and find y.